Find the slope of any line perpendicular to the line through points (0,5) and
(-3,-4).

Answer by checkley75(3666)

SLOPE OF A LINE THROUGH (0,5) & (-3,-4) IS:
(-4-5)/(0+3)=-9/3=-3 THEREFORE THE SLOPE FOR ANY LINE PERPENDICULAR TO THIS LINE IS -1/3. THEN THE EQUATION OF ANY PERPENDICULAR IS:
Y=-X/3+b